OpenStudy (1018):

for y=4sin(3x), find y(pi)

OpenStudy (welshfella):

plug in pi for x

OpenStudy (1018):

oh wait sorry that's supposed to be y'(pi)

OpenStudy (welshfella):

ok so you have to find the derivative of 4 sin3x

OpenStudy (1018):

oh wait i think i got it can you check if this is right

OpenStudy (welshfella):

ok

OpenStudy (1018):

y'=4cos(3x)(3) ?

OpenStudy (welshfella):

yea or you could write it as 12 cos (3x)

OpenStudy (1018):

thanks! may i ask, is this always the right way, because at first i thought i would use the chain rule on 4 and sin(3x)

OpenStudy (welshfella):

sin (3x) is a compound function Ignore the 4 at first - just use chain rule on sin 3x . Finally multiply the result by 4.

OpenStudy (1018):

oh, so should i just ignore numbers that appears before the trig function? i should just do what's inside the funtion?

OpenStudy (welshfella):

yes - but of course dont forget to operate on them after you have done the differentiation

OpenStudy (welshfella):

eg 2 cos 3x = 2 * -3sin 3x = -6 sin 3x

OpenStudy (1018):

yes, thanks! oh wait, last one, may you answer this? i think the answer is 0 but not sure

OpenStudy (1018):

y=1-2u u = x^3 - 2x^2 + 6 find y''(x=1)

OpenStudy (welshfella):

y = 1 - 2(x^3 - 2x^2 + 6) y = 1 - 2x^3 + 4x^2 - 12 what do you get for y' ?